OpenStudy (anonymous):

David opens a bank account with $140 and then deposits $150 each month. Louis opens a bank account with $260 and deposits $120 each month. How many months will it take for David and Louis to have equal amounts of money in their accounts? A. 2 B. 4 C. 5 D. 7

OpenStudy (anonymous):

plz help Medal to best answer and Fan

OpenStudy (anonymous):

@Michele_Laino plz help

OpenStudy (perl):

you have two linear equations

OpenStudy (perl):

the rate of the first account is 150 per month, and it starts at 140 y = 150*x + 140 the rate of the second account is 120 per month and starts at 260 y = 120*x + 260 Now where do these two lines intersect

OpenStudy (michele_laino):

we can mathematize both bank accounts, with a linear function, as @perl well said

OpenStudy (anonymous):

120+260=380

OpenStudy (michele_laino):

please you have to solve the system of the two equations that @perl wrote above

OpenStudy (anonymous):

i don't get

OpenStudy (anonymous):

i don't get it

OpenStudy (michele_laino):

first equation: \[y=140+150x\]

OpenStudy (michele_laino):

second equation: \[y=260+120x\]

OpenStudy (anonymous):

290

OpenStudy (michele_laino):

please you have to find x such that both bank accounts are equal, so: \[260+120x=140+150x\] please, solve for x

OpenStudy (anonymous):

i am going to do a different test i don't get anything you are telling me so sorry

OpenStudy (anonymous):

i guessed and got 80% and passed i said it was B

OpenStudy (michele_laino):

I solve the equation above!

OpenStudy (michele_laino):

we have: \[30x=120\] from which: \[x=4\] so after 4 months david and Louis have the same quantity of money

OpenStudy (michele_laino):

x is number of months y is the money quantity, so after 4 months, both David and Louis have 740$

OpenStudy (michele_laino):

since 150*4+140=740, and 120*4+260=740

OpenStudy (michele_laino):

is it all clear, please? @TeenWolfGirl

OpenStudy (michele_laino):

so B.

OpenStudy (anonymous):

yes thank you

OpenStudy (michele_laino):

thank you! @TeenWolfGirl